# An Empirical Study on LLM-based Agents for Automated Bug Fixing

### v1

@@ 4.2. RQ2: Effectiveness of FL
Fault localization is a critical step in resolving issues; only when the faulty code element is accurately identified can the model generate a semantically correct patch. In this RQ, we examine the effectiveness of seven methods in fault localization. As some tools’ run trajectories lack explicit fault localization results or provide insufficient localization granularity—offering bug location information only at the file level, without finer-grained function or line-level details—we use the modified files and line information in the final submitted patch as a reference to fairly assess each tool’s localization capability. Specifically, we downloaded the final patch set submitted for each tool’s benchmark on the SWE-bench Lite website, then parsed the formatted patch files. From the perspective of the files pre-modification (pre-patch code files), we recorded the files and lines that were modified. We applied the same processing to the manually crafted patches (ground truth).
@@
After obtaining the above processed data, we created Figure 1(b) and Figure 1(c) to illustrate the hit rate of fault localization at the file and line levels, respectively. The meaning of the different modules in the figures is similar to Figure 1(a), except that the latter is oriented towards the issue resolve task. For the file-level localization task, since cases in SWE-bench Lite only modify one file (the ground truth file), but repair tools are not limited to a single file in generating patches, we consider it a file hit if the patch file set contains the ground truth file. For line-level localization, we define a hit as an overlap between the lines modified in the patch and the ground truth line.
Based on this definition of a hit, we investigated the localization performance of the seven tools. As shown in Figure 1(a), MarsCode Agent performs best on the file-level localization task, successfully localizing 239 cases. It is followed by Alibaba Lingma Agent, Gru, Agentless+RepoGraph, AutoCodeRover, Honeycomb, and finally Agentless, which successfully localized 206 cases. MarsCode Agent shows a 16.0% improvement over Agentless. As the highest and lowest-ranking tools in the issue resolve task, MarsCode Agent and Agentless demonstrate a smaller performance gap of 43.9% in the issue resolve task, which is much larger than their difference in file-level localization performance. Additionally, Agentless+RepoGraph, another representative of non-agent systems, ranks fourth in file-level fault localization but sixth in the issue resolve task. These comparisons indicate that the file-level localization task is generally simpler than the subsequent patch generation task. Moreover, agent-based repair methods like MarsCode Agent show more strength in subsequent patch generation optimization.
However, we encountered an unexpected result: Honeycomb, the second-best tool in the issue resolve task (with minimal gap from MarsCode), performs poorly in file-level localization, ranking sixth, with only two more hits than the lowest-ranking Agentless. To investigate this further, we examined the tools’ performance in line-level localization. As shown in Figure 1(c), MarsCode remains the best method, while Honeycomb ranks a close second, trailing by only two hits. This suggests that although Honeycomb performs poorly in file-level localization, it has a relatively high hit rate for faulty lines within the located file (137/208 = 65.9%). Therefore, strong line-level localization performance is the main reason Honeycomb resolves a larger number of issues. This further suggests that line-level localization results are more closely associated with the subsequent patch generation process than file-level results. Improving fine-grained line-level localization may thus be key to enhancing end-to-end repair effectiveness, aligning with our conclusions based on Table 3 (i.e., the line-level location metric shows the most significant difference between the All-resolve Set and No-one-resolve Set).
Overall, three key insights can be derived from the above analysis: (1) Existing methods have achieved relatively strong performance in file-level fault localization, but there remains room for improvement in line-level localization tasks. (2) Finer-grained line-level localization results are more closely correlated with end-to-end issue resolution. When erroneous code lines can be accurately identified by a tool, the generation of a correct patch becomes more feasible.
@@
Insights(9)-(10): Overall, two key insights can be derived from the above analysis: (9) Existing methods have achieved relatively strong performance in file-level fault localization, but there remains room for improvement in line-level localization tasks. (10) Finer-grained line-level localization results are more closely correlated with end-to-end issue resolution. When erroneous code lines can be accurately identified by a tool, the generation of a correct patch becomes more feasible.
4.3. RQ3: Effectiveness of Reproduction
Table 5. Reproduction Statistics of Leading LLM-based Bug Fixing Systems on SWE-Bench Lite. #Resolved refers to the number of successfully resolved issues, #(Reproduced&Resolved) refers to the number of issues that were successfully resolved and involved bug reproduction during the resolution. %(Reproduced&Resolved) refers to #(Reproduced&Resolved) / #Resolved.
System Name Type Reproducer #Resolved #(Reproduced&Resolved) %(Reproduced&Resolved)
MarsCode Agent Commercial 118 83 70.3
Honeycomb 115 115 100
Gru $\times$ 107 0 0
Alibaba Lingma Agent 99 12 12.1
AutoCodeRover Open Source 92 39 42.4
Agentless + RepoGraph $\times$ 89 0 0
Agentless $\times$ 82 0 0
We analyzed the trajectories of the seven systems evaluated on SWE-Bench-Lite, examining the use of reproduction in different systems and analyzing some typical cases to gain a deeper understanding of the impact of reproduction on bug fix success rates, as well as the challenges reproduction faces in current systems. Table 5 presents the statistics on reproduction usage across different systems. The systems Gru, Agentless, and Agentless+RepoGraph do not include a reproduction module, while Marscode Agent, Honeycomb, Alibaba Lingma Agent, and AutoCodeRover all employ a reproduction module, and in the successfully fixed bugs, they perform bug reproduction for 70.3%, 100%, 12.1%, and 42.4% of the issues, respectively. Marscode Agent, Alibaba Lingma Agent, and AutoCodeRover decide whether to perform bug reproduction based on the issue, opting for static analysis for bugs that are difficult to reproduce. Honeycomb, on the other hand, requires the LLM to attempt bug reproduction for all issues.
Through an analysis and comparison of the trajectories of different systems, we found that reproduction can provide additional information for defect localization when issue information is lacking, and it helps verify the accuracy of generated candidate patches. Our trajectory analysis revealed that among the 168 issues resolved by at least one system, 24 issues could only be solved using reproduction. These 24 instances generally had brief textual descriptions that only briefly outlined the bug conditions, making defect localization challenging based on the text alone. Of these, 20 instances included code snippets in the issue description to help understand the bug. For example, in sympy-15346, the issue contains only a brief textual description followed by an example of the bug. Locating the defect based solely on the text was challenging. Marscode Agent implemented a bug reproduction script based on the example in the issue and gradually identified the defect location based on the reproduction results, eventually generating the correct patch. Notably, during trajectory analysis, we observed that Marscode Agent initially struggled to pinpoint the defect location when running the reproduction script. It then attempted to print intermediate variables at different points in the script, and it successfully identified the defect location by analyzing the outputs during execution. This further demonstrates the advantage of the autonomy of agent-based systems in bug fixing.
Although reproduction could provide additional information and verify the correctness of candidate patches, it does not always improve the success rate of bug fixing. For instance, in django-11422, the "Autoreloader" mentioned in the issue description directly points to the file $autoreload.py$. By using information such as "track changes" from the issue, Agentless accurately located the code snippet needing modification and generated the correct patch. However, systems like Marscode Agent and Honeycomb, which performed bug reproduction, was distracted by the reproduction information. The reproduction diverted the LLM’s attention away from the issue description, leading to incorrect localization of the defect cause and ultimately failing to produce the correct patch.
Insights(11)-(12): (11) Reproduction can provide additional information for defect localization when issue information is lacking and help verify the accuracy of generated candidate patches. (12) When the issue description is already clear and precise, reproduction may mislead the LLM’s judgment, reducing its focus on the issue description.
@@ 5.2. Agentic Flow
From the Agentic flow perspective, agents should especially focus on the quality of the issue and pay attention to multiple suspicious locations in the stack trace. The Agentic flow design should include mechanisms to check the completeness of patches and consider the global impact of the fixes. During the use of the model, mechanisms should be established to either avoid the randomness of the model’s output or make full use of the diversity in the model’s output.
In fault localization, the accuracy of line-level localization is more important than file-level, as the discovery space at the line level is larger, necessitating finer-grained localization results. During the reproduction process, it is crucial to strengthen the determination of the correctness of the reproduction, as an incorrect reproduction can lead to the failure of the entire solving process.
@@ 6. Threats to Validity
Uncertainty of LLM: The output of LLMs is stochastic, leading to a probabilistic nature for whether an instance is solved. In this work, we directly analyzed the patches and trajectories submitted by various systems, assuming that the results submitted to SWE-bench represent the best performance of the Agent. Furthermore, conducting multiple experiments for each system to eliminate stochasticity is impractical in terms of both cost and accessibility.
@@ 7. Related Work
In this section, we discuss basic concepts of large language models and their application on software engineering tasks, especially for fault localization and automated program repair. We also discuss recent advances in LLM-based agents for software engineering.
@@ 7.1. Large Language Models
Large language models (LLMs) are highly advanced pre-trained language models. These models undergo initial unsupervised training on vast amounts of corpus, followed by fine-tuning for specific tasks to enhance performance. In natural language processing (NLP), LLMs have been extensively applied to various tasks such as machine translation (Wang et al., 2023d; Zhang et al., 2023b), text summarization (Zhang et al., 2023c), and classification (Mayer et al., 2023).
@@
Large language models (LLMs) have become a promising choice for various software engineering tasks due to their impressive performance in both code generation and understanding (Yang et al., 2023). Researchers and developers have applied LLMs to several software engineering tasks, such as program synthesis (Liu et al., 2024b; Zhu-Tian et al., 2024; Wang et al., 2023b, c, a), code translation (Yu et al., 2024; Yang et al., 2024b), program repair (Lin et al., 2024; Jiang et al., 2023; Xia et al., 2023), fault detection and localization (Du et al., 2024; Qin et al., 2024), incident analysis (Chen et al., 2024; Ahmed et al., 2023), code summarization (Geng et al., 2024) and testing (Sun et al., 2023). For example, Codex (Chen et al., 2021), StarCoder (Lozhkov et al., 2024), and DeepSeek-Coder (Zhu et al., 2024) are notable code-specific LLMs developed through extensive training on large datasets of open-source code snippets. Additionally, instruction-following code-specific LLMs such as DeepSeek-Coder-Instruct (Zhu et al., 2024) and Magicoder (Wei et al., 2023) have been created using instruction-tuning methods to enhance their utility in coding tasks.
7.2. Fault Localization

### v2

@@ 4.2. RQ2: Effectiveness of FL
Fault localization (FL) is a critical step in resolving issues, and only when the faulty code element is accurately identified can the model generate a semantically correct patch. In this RQ, we examine the effectiveness of six methods in FL. As run trajectories of most repair techniques lack intermediate FL results or provide insufficient localization granularity (e.g., only offering file-level FL information, without finer-grained code symbol-level or line-level details), we extract the localization information in different levels from the final submitted patch to fairly assess each tool’s localization capability. Specifically, we downloaded the final patch set submitted by each tool on the official SWE-bench Verified website, then parsed the formatted patch files. From the perspective of the files pre-modification (pre-patch code files), we recorded the files and code symbols that were modified. We did not conduct line-level experiments because we believe that individual lines of code do not adequately represent a complete functional module, and therefore, have limited reference value. We applied the same processing to the golden patches.
@@
After obtaining the above processed data, we created Figure 2 and Figure 3 to illustrate the hit rate of FL at the file and code symbol levels, respectively. The meaning of the different modules in the figures is similar to Figure 1, except that Figure 3 is oriented towards the FL task.
Furthermore, we analyzed the localization performance of the six tools based on their definition of a hit. As shown in Figure 2, the W&B Programmer outperformed others in the file-level localization task, successfully localizing at least one buggy file in 448 cases (Figure 2(a)), and all buggy files in 394 cases (Figure 2(b)). In the "Cover-at-least-one-file" setup, W&B Programmer is followed by Devlo, Blackbox AI Agent, Emergent E1, Midwit Agent, and finally Learn-by-interact with 342 cases localized. This indicates a 31.0% improvement for W&B Programmer over Learn-by-interact. In the "Cover-all-files" scenario, only Devlo and Blackbox AI Agent swapped their positions, while the rest remained in the same order. The top-ranked tool, W&B Programmer, showed a 25.5% improvement over the lowest-ranked Learn-by-interact.
Additionally, since each instance may have multiple buggy files and each system may predict a list of files rather than a single file, we compiled the prediction accuracy, recall, and F1-score for each system at the code file level. This data is provided after the name of each system in Figure 2. Based on the F1-score, the systems, sorted from lowest to highest, are Learn-by-interact, Emergent E1, Devlo, Midwit Agent, Blackbox AI Agent, and W&B Programmer. Excluding Learn-by-interact, the relative order of the remaining systems aligns with their ranking in the issue resolution task, indicating a strong positive correlation between accurate file-level localization and overall bug-fixing performance.
Next, we assess FL performance at the code symbol level. In the "Cover-at-least-one-symbol" scenario, the top bug-fixing system is W&B Programmer, successfully localizing 396 cases. The lowest bug-fixing system is Learn-by-interact, with 308 cases localized, resulting in a 28.6% improvement for W&B Programmer. In the "Cover-all-symbols" scenario, W&B Programmer localized 50 more cases than Learn-by-interact, showing a 20% improvement. Compared to file-level results, code symbol-level localization shows more variance from the issue resolution task, and the abilities of various indicators differ. For example, W&B Programmer ranks highest in F1-score for code symbol prediction, making it an all-around performer, while Midwit Agent excels in prediction precision. Although Blackbox AI Agent does not rank in the top three for precision and recall, it remains the second-best in localization from an instance perspective.
We believe these differences are primarily due to the varying localization strategies at the code symbol granularity employed by different models. Some repair systems might focus on recalling more code symbols to provide more flexibility for subsequent patch generation. Others might aim to precisely identify the first erroneous code symbols to avoid misguidance in subsequent repair processes, without overly considering the completeness of code symbol recall and associated modifications. Despite these differences, we still observe a positive correlation between code symbol-level FL results and bug-fixing results. Additionally, compared to the file-level, there is still considerable room for improvement in symbol-level localization effectiveness in the future.
It is easy to observe that in localization experiments, Learn-by-interact consistently ranks last and performs significantly below the other five repair systems. However, in the bug-fixing task, it ranks fourth. To investigate this, we constructed three case sets: 301 cases that Learn-by-interact successfully repaired (Set-Repair), 342 cases where it successfully localized at least one erroneous file (Set-File-Localization), and 308 cases where it localized at least one erroneous code symbol (Set-Symbol-Localization).
[…]
In comparison, these probabilities for W&B Programmer are 69.6% and 72.7%, respectively. Thus, while Learn-by-interact’s success rate in predicting the correct files and code symbols is minimal, once a prediction is successful, the subsequent repair success rate is remarkably high. This is why Learn-by-interact can achieve relatively good repair outcomes despite its poor FL performance.
@@
Insights: (7) Existing methods have achieved relatively strong performance in file-level FL, but there remains room for improvement in code symbol-level localization tasks. (8) Both coarse-grained file-level FL results and fine-grained code symbol-level FL results are positively correlated with final bug-fixing performance.
4.3. RQ3: Effectiveness of Bug Reproduction
Under most of top-leading bug-fixing systems, the core process involves searching for suspicious codes, generating bug reproduction scripts, successfully reproducing the issue, and modifying the functional code until the issue is resolved (i.e., passing the reproduction testings). The generation of bug reproduction scripts is of paramount importance for the model’s understanding of the problem and for verifying the correctness of the generated patches. Therefore, in this RQ, we plan to specifically investigate the significance of this task. An initial idea was to extract the reproduction scripts generated during the repair process from the trajectories submitted by the top six repair methods. However, we encountered numerous difficulties during the attempts.
Reproduction scripts within a project have variable names and locations, which complicates the accurate extraction of filenames using rule-based methods. These scripts may be named differently and stored in various directories, adding complexity. Additionally, trajectories do not directly provide the final versions of reproduction files, as repair methods often modify the initially visible contents using code editing tools. These modifications are partially visible, yet the full text remains obscured. The proprietary nature and diverse designs of these tools make understanding and fitting their functions challenging. Moreover, the execution commands for these files also vary. Some methods generate new Python files with a main entry point, while others alter existing test files and execute them using frameworks such as pytest. Extracting these commands from trajectories is thus a difficult task. Finally, problem reproduction processes may involve altering the running environment, such as installing dependencies via pip or exporting environment variables, which, if not recorded, can impede the execution of reproduction scripts effectively.
Having abandoned the idea of extracting and analyzing reproduction scripts from existing repair systems, we considered implementing an Agent with a standard process to preserve the final reproduced script during its runtime, thereby advancing subsequent investigations. Specifically, we designed a repair framework with two Agents, named RepoFixer. Compared with other SOTA bug-fixing systems, for example, Agentless employs a single-step, phase-separated inference process, while SWE-agent uses a multi-step ReAct-based framework. Our implementation aligns more closely with the latter, but simplifies the agent architecture to retain only core components—avoiding heavy optimizations—to better generalize insights for bug reproduction tasks. In RepoFixer, the first Agent, called Searcher, takes as input the issue description from each sample in the SWE-bench Verified dataset and the current commit state of the code repository. We instruct the Searcher via system prompt commands to perform a code symbol granularity FL task and, through multi-turn conversations, continuously search and read the context related to the issue description, ultimately returning a list of suspicious code symbols.
Our second Agent, called Fixer, takes as input the issue description for each sample and the list of suspicious code symbols returned by the Searcher, aiming to generate a patch that resolves the issue. In the system prompt, we suggest the Fixer’s workflow: first, generate a problem reproduction script, then run it and self-assess whether the issue is successfully reproduced based on the run results. If the issue is not successfully reproduced, continue modifying the reproduction script until it is deemed that the issue described in the issue has been successfully reproduced. Then, generate the patch and use the reproduction script to verify whether the modified code successfully resolves the issue (i.e., the previously reproducible issue no longer occurs). If successful, save the patch and end the process; otherwise, continue to repair the patch and run the reproduction script until the issue is considered resolved or the predefined number of dialogue turns is reached. Regardless of whether a patch is successfully generated, we will save the latest edited reproduction script and the corresponding execution commands.
During this process, we utilized the powerful Claude 3.5 Sonnet model (October 2024 version) [^7] https://www.anthropic.com/news/3-5-models-and-computer-use as the base model to implement the Agent process. We used Claude’s official toolkits, namely bash and str_replace_editor, as the Agent runtime toolkits and utilized the standard Function Call interface for model requests. Post-experimental statistics show that our RepoFixer successfully resolved a total of 255 issues, accounting for 51% of all samples. While we did not adopt very heavy sampling or regression testing strategies, and complex mechanisms such as self-reflection, it appears that there remains some gap in performance compared to the top-ranked methods. However, the current focus of our research question is on investigating the quality of reproduction scripts generated by Agent systems. To explore this, we designed the following experiment.
We placed the collected reproduction scripts back into the Docker environment for each sample and then ran the script once in the repository before applying the golden patch, recording the output as before_fix_status. Subsequently, in a new Docker instance, we applied the golden patch to the code repository and then ran the same reproduction script again, recording the output as after_fix_status. The golden patch refers to the correct repair patches provided in the SWE-bench Verified dataset. A natural consideration is that if the output of the same reproduction script changes before and after applying the golden patch, it can be assumed that the reproduction script is related to the issue that the golden patch addresses. Thus, reproduction scripts producing such status differences should be considered of higher quality, whereas those without status differences seem likely to have a weaker association with the issue.
Based on this, we collected experimental results. Out of 500 instances generating 500 reproduction scripts, 384 scripts produced status differences before and after the golden patch application to the repository, representing 76.8%. We term reproduction scripts exhibiting this property as "related reproduction scripts." We hypothesize that the higher the quality of the reproduction information in the issue description, the more likely the model is to generate related reproduction scripts. For this, we conducted detailed statistics. Among 194 samples containing "Contains REs," 156 produced related reproduction scripts, accounting for 80.4%; among 169 samples containing "Contains Partial REs," 135 produced related reproduction scripts, accounting for 79.9%; among 87 samples where the information is in natural language ("Info in NL"), 61 produced related reproduction scripts, accounting for 70.1%; and among 50 samples marked "Not Enough Info," 32 produced related reproduction scripts, accounting for 64%. It can be observed that as the quality of reproduction descriptions in the issue decreases, the proportion of generating related reproduction scripts also decreases. Thus, for users, accurately and comprehensively recording the process of triggering an issue in code form is of significant value for automating the resolution of the issue.
Insights: (9) The most effective bug-fixing systems currently employ a process based on bug reproduction, patch generation, and patch correctness verification. Among these, the bug reproduction process holds significant value. (10) Drafting clear and comprehensive bug reproduction cases within the issue can greatly reduce the difficulty of reproducing the bug, consequently enhancing the efficiency of end-to-end bug-fixing tasks.
@@ 5.2. Agentic Flow
Agents should especially focus on the quality of the issue and pay attention to multiple suspicious locations in the stack trace. The Agentic flow design should include mechanisms to check the completeness of patches and consider the global impact of the fixes. During the use of the model, mechanisms should be established to either avoid the randomness of the model’s output or make full use of the diversity in the model’s output. The effectiveness of fault localization is important for bug-fixing tasks, finer-grained code symbol level fault localization task still has significant room for improvement. During the reproduction process, it is necessary to increase the proportion of the dynamic repair process, centered on bug reproduction, patch generation, and patch validation. The quality of bug reproduction scripts should be improved by introducing more relevant artifacts and increasing model reflection.
5.3. General Applicability
While our study uses the SWE-bench Verified benchmark with only Python repositories, we believe the analytical insights generalize beyond this setting. The benchmark is widely used by organizations such as OpenAI and Anthropic, and the analyzed systems rely on language-agnostic tools, with no signs of Python-specific tuning. Nevertheless, we recognize the importance of multilingual benchmarks and are developing a broader dataset for future evaluations.
5.4. Qualitative Evaluations from Developers
While our evaluated patch correctness via test suite validation, we acknowledge it may not fully capture patch quality from a developer’s perspective. To supplement this, we manually analyzed 30 patches per system, evaluating code style, readability, and adherence to development norms. We found that agent-based workflows enable context-aware understanding, produce syntactically correct code, reuse existing structures effectively, and often yield clean, human-readable patches with meaningful comments
These results indicate that modern LLM-based agents can generate patches that are both correct and stylistically aligned with professional standards. A more comprehensive evaluation with developer feedback remains future work.
@@ 6. Threats to Validity
Uncertainty of LLM: The output of LLMs is stochastic, leading to a probabilistic nature for whether an instance is solved. In this work, we directly analyzed the patches and trajectories submitted by various systems, assuming that the results submitted to SWE-bench represent the best performance of the Agent. Furthermore, conducting multiple experiments for each system to eliminate stochasticity is impractical in terms of both cost and accessibility. As these systems are typically closed-source, we cannot standardize their model and decoding configurations or conduct a controlled ablation using identical LLM backbones. Thus, differences in model capacity inherently contribute to variations in performance, a factor we acknowledge explicitly as a limitation of our comparative analysis.
LLM-based Issue Quality Assessment: To ensure that the LLM’s evaluation results are as close as possible to those of software experts, we took the following measures: (1) We selected the DeepSeek-R1 model, which has top-tier inference performance, as the evaluator; (2) Initially, we conducted model annotations and manual annotations on a subset of the SWE-bench Verified dataset, continuously refining the scoring prompts based on discrepancies until they aligned with human expectations; (3) We used DeepSeek-R1 to perform five annotations for each metric of each instance and selected the option with the highest frequency as the final choice.
Reliability of Reproduction Scripts: Another potential threat to validity lies in the possibility of false positives—scripts marked as “related” that do not actually trigger the bug. While our approach leverages output divergence as an indication of successful bug reproduction, inspired by mutation-based fault localization (Li and Zhang, 2017), this method may still misclassify some non-triggering scripts as successful reproductions. Incorporating additional checks, such as verifying log patterns or exception types, could theoretically reduce such misclassification. However, given the practical complexity of implementing and reliably automating these checks within limited agent interactions, we consciously accepted this limitation to maintain a balance between practicality and effectiveness. Future work could explore incorporating such checks to more precisely measure reproduction accuracy and further mitigate this validity threat.
@@ 7.1. Large Language Models
Large language models (LLMs) are highly advanced pre-trained language models. These models undergo initial unsupervised training on vast amounts of corpus, followed by fine-tuning for specific tasks to enhance performance.
@@
Large language models (LLMs) have become a promising choice for various software engineering tasks due to their impressive performance in both code generation and understanding (Yang et al., 2023), such as program synthesis (Liu et al., 2024b; Zhu-Tian et al., 2024; Wang et al., 2023b, c, a), code translation (Yu et al., 2024; Yang et al., 2024b), program repair (Lin et al., 2024; Jiang et al., 2023; Xia et al., 2023), fault detection and localization (Du et al., 2024; Qin et al., 2024), incident analysis (Chen et al., 2024; Ahmed et al., 2023), code summarization (Geng et al., 2024) and testing (Sun et al., 2023). For example, Codex (Chen et al., 2021), StarCoder (Lozhkov et al., 2024), and DeepSeek-Coder (Zhu et al., 2024) are notable code-specific LLMs developed through extensive training on large datasets of open-source code snippets. Additionally, instruction-following code-specific LLMs such as DeepSeek-Coder-Instruct (Zhu et al., 2024) and Magicoder (Wei et al., 2023) have been created using instruction-tuning methods to enhance their utility in coding tasks.
7.2. Fault Localization and Program Repair